It is known by Trade Kings that in the manufacturing of Chelsea Biscuits, fixed costs per month are K15,000. If the marginal cost is F(Q) = 0.000204Q2 - 0.046Q + 6 when Q = 25 . Find the total cost function. 


1
Expert's answer
2021-09-12T23:51:51-0400

Given

"dC\/dQ=F(Q)=0.000204Q^2-0.046Q+6"

Then


"C=\int(0.000204Q^2-0.046Q+6)dQ"

"=0.000068Q^3-0.023Q^2+6Q+K"

Given "K=15000" Therefore


"C(Q)=0.000068Q^3-0.023Q^2+6Q+15000"

If "Q=25"


"C(25)=0.000068(25)^3-0.023(25)^2+6(25)+15000"

"=15136.69"